Which of the following equations represent the product of two consecutive odd
integers that equal 63?
A x(x − 1) = 63
B x(x + 1) = 63
C x(x − 2) = 63
D x(x + 2) = 63
E (x + 9)(x − 7) = 0
F (x + 9)(x + 7) = 0

Found 5 solutions by mananth, Alan3354, josgarithmetic, ikleyn, greenestamps:
Answer by mananth(16946)

(x + 9)(x + 7) = 0

roots are -9,-7
They are consecutive odd integers

Answer by Alan3354(69443)

Which of the following equations represent the product of two consecutive odd
integers that equal 63?
A x(x − 1) = 63
B x(x + 1) = 63
C x(x − 2) = 63
D x(x + 2) = 63
E (x + 9)(x − 7) = 0
F (x + 9)(x + 7) = 0
===============
C, D & F
----------------
F (x + 9)(x + 7) = 0
@ x = 0, 7*9 = 63

Answer by josgarithmetic(39618)
Answer by ikleyn(52788)

.

This question is a mathematical joke problem.

Its goal is to check / (to test) if a person is able to read a simple mathematical text
and if he  (or she)  does really understand the meaning of the terms that he  (or she)  reads.

The answer is :   equations  (C)  and  (D).


Those who can answer correctly go to the left; the rest go to the right.



Answer by greenestamps(13200)



Equations C and D are both equations that represent the statement "the product of two consecutive odd numbers is 63."

Equation F has solutions that are two consecutive odd numbers whose sum is 63; but that equation does NOT represent the statement that the product of two consecutive odd numbers is 63.

ANSWERS: C and D (only)